### v1

# Fractal Patterns in Discrete Laplacians: Iterative Construction on 2D Square Lattices

Abstract
We investigate the iterative construction of discrete Laplacians on 2D square lattices, revealing emergent fractal-like patterns shaped by modular arithmetic. While classical $2$-style iterations reproduce known structures such as the Sierpiński triangle, our alternating binary–ternary (2322-style) process produces a novel class of aperiodic figures. These display low density variance, minimal connectivity loss, and non-repetitive organization reminiscent of Dekking’s sequences. Fourier and autocorrelation analyses confirm their quasi-periodic nature, suggesting applications in self-assembly, sensor networks, and biological modeling. The findings open new paths toward structured randomness and fractal dynamics in discrete systems.
These findings also open avenues for exploring higher-dimensional Laplacian constructions and their implications in quasicrystals, aperiodic tilings, and stochastic processes.
keywords:
discrete Laplacian, modular arithmetic, fractals, aperiodic patterns, 2D lattices, Sierpiński triangle, Dekking sequence, sensor networks
journal: Fractals
[…]
organization=Faculty of Natural and Technical Sciences, The John Paul II Catholic University of Lublin,addressline=Konstantynow 1H, city=Lublin, postcode=20-708, country=Poland, email: malnow@kul.pl
{graphicalabstract}[Uncaptioned image]
Introduction
The discrete Laplacian is a fundamental mathematical operator with applications in physics, biology, and computational science. It plays a crucial role in graph theory, numerical analysis, and dynamical systems, serving as a discrete analog to the continuous Laplace operator. Iterating the discrete Laplacian on 2D square lattices gives rise to complex emergent structures, often displaying fractal-like organization.
Previous studies of Discrete Laplacians on 2D lattices have documented visually striking fractal patterns, such as snowflake-like and carpet-like motifs, arising from binary iterations alone Aiba et al. (2006). These patterns emerge due to modular arithmetic constraints and exhibit diverse forms, depending on the chosen seed and neighborhood structure. However, binary (modulo 2) iterations tend to produce periodic structures and, at specific iterations, dissociate into their initial seed configurations. This motivates the exploration of higher-order modular arithmetic, such as ternary (modulo 3) and quaternary (modulo 4) iterations, which significantly expand the space of possible patterns and often result in aperiodic, structurally richer formations.
In this paper, we investigate a special class of iterative Laplacian constructions, focusing on the interplay between binary and ternary arithmetic in pattern formation. We introduce the 2322-style pattern, an alternating sequence of binary and ternary iterations, which gives rise to structurally distinct, non-repetitive figures. Unlike purely binary $2$-style figures, which exhibit dissociative periodicity, 2322-style figures display low density variance, minimal connectivity loss, and quasi-aperiodicity. These properties are reminiscent of Dekking’s non-repetitive sequences, which play a significant role in combinatorial mathematics and aperiodic order.
Beyond their mathematical significance, these structures may have applications in physics, materials science, and computational modeling. The concluding section of this paper will discuss their potential relevance, including connections to biological self-organization, sensor technology, material science, and fractal-based computational algorithms.
The paper is structured as follows: Section 1 introduces the iterative construction process; Section 2 explores binary figures and Sierpiński-like patterns; Section 3 presents the distinguishing properties of 2322-style figures; Section 4 compares cell sequence statistics, including Fourier and autocorrelation analysis; and Section 5 discusses applications and outlines directions for future research.
1 Iterative Dynamical Systems of Discrete Laplacians
1.1 Initial Conditions and Seed Configurations
We consider an automaton initialized with binary values:
$\nu_{0}(p)=\begin{cases}1,\quad&\text{if }p\text{ belongs to the seed,}\\ 0,\quad&\text{otherwise}.\end{cases}$
where 1 represents an occupied cell ( ) and 0 denotes an unoccupied cell ($\square$).
The choice of seed configuration significantly influences the evolutionary dynamics of the system. Some commonly used seeds include:
  • 1.
    Single point (minimal seed, simplest growth),
  • 2.
    Line segment (introducing directional spread),
  • 3.
    Geometric clusters (leading to more intricate self-organization).
Figure 1: Examples of seed configurations.
1.2 Automaton Evolution
The iterative update rule follows the form:
$\nu_{i}(p)=\sum_{g\in Ne(p)}\big{(}\nu(g)-\nu(p)\big{)}\mod n,$ (1)
where:
  • 1.
    $n$ is the modulus (defining binary, ternary, or higher-order arithmetic),
  • 2.
    $i$ represents the iteration step,
  • 3.
    $Ne(p)$ defines the neighborhood structure, governing local interactions.
Figure 2: Examples of neighborhoods.
1.3 Modular Arithmetic and Iterative Extensions
To extend beyond binary figures, we introduce modular arithmetic with $n>2$. Each color represents a distinct modular class:
  • 1.
    Binary iteration ($n=2$): classical black-and-white figures.
  • 2.
    Ternary iteration ($n=3$): introduces a new state, increasing complexity.
  • 3.
    Quaternary and higher ($n>3$): potential multi-state structures.
We define a $k$-nary iteration:
$u_{i}(p)=\sum_{g\in Ne(p)}\big{(}u(g)-u(p)\big{)}\mod k,\quad k=2,3,4,\dots$
where the iteration is binary for $k=2$, ternary for $k=3$, and so forth.
In this paper, we examine specific iteration sequences:
$u_{i}(p)=\begin{cases}\sum\limits_{g\in Ne(p)}\big{(}u(g)-u(p)\big{)}\mod n,&\text{if }i\equiv 2\pmod{4},\\ \sum\limits_{g\in Ne(p)}\big{(}u(g)-u(p)\big{)}\mod 2,&\text{otherwise}.\end{cases}$ (2)
This generates $222n$ 22-style figures, including:
  • 1.
    $2$-style (purely binary evolution),
  • 2.
    2322-style (introducing ternary interference in specific steps).
Figure 3: Binary iterations from a Sierpiński seed and Neumann neighborhood.
2 Binary style figures and their properties
Prior work on binary style figures by Suzuki and Maegaito uncovered visually striking configurations, including snowflake-like, butterfly-like, and Persian carpet-like motifs Aiba et al. (2006). Theoretical developments, including fixed-point theorems, periodicity results, and connections to binomial and trinomial sequences have been found Hadlich et al. (2011).
We shall investigate binary style figures via many seeds and neighbourhoods. First we shall describe characteristic feature of binary figures and the suprising appearance of Sierpinski-like triangles.
To analyze connectedness, we shall say the figure is $k$-steps away from connectedness, $k>0$, $k\in\mathbb{Z}$, if, in order to make it connected, it is required to add paths between its connected components, the longest being of length $k$.
A connected figure ($A$), figures: 1-step away (figure $B$), and 2-steps away from connectedness (figure $C$) are shown below.
All binary-style figures posses a special feature which is described in:
Proposition. Every binary-style figure of arbitrary seed and neighborhood at iterations $i=8k$, $k=1,2,3,\ldots$ consists of a spread of its seeds and is at least 13-steps away from connectedness.
An illustration of this proposition is shown in Fig. 3. The sketch of the proof will be given in the Appendix.
Generalization to Other Neighborhoods
Although this proof is carried out for Diag-Neumann, all observed figures for various seeds and neighborhoods exhibited the same behavior. Therefore, it is highly likely that all $222n$ 22-style figures for even $n$ behave similarly, all the figures are reduced to a spread of seeds at iterations $8k$.
2.1 Density Measure
The density of a figure provides a quantitative measure of its occupancy on the lattice and offers insights into its connectivity properties. A low density often correlates with fragmentation, where the figure consists of multiple disconnected components.
Definition. The density $\rho(i)$ of a figure at iteration $i$ is defined as the ratio of occupied cells to the total available lattice area:
$\rho(i)=\frac{\sum_{p\in\text{lattice}}\text{sgn}(f(p))}{(3+2i)^{2}}.$
Since binary figures are observed to become a spread of seeds at iterations $i=8k,\,k=1,2,\dots$, their density exhibits sharp reductions at these steps (see Fig. 8 (a)). This pattern is consistent with our analytical findings, suggesting that binary figures experience periodic dissociation into separated copies of their initial seeds.
At iteration $i=8$, density does not exceed:
$\rho(8)\leq\frac{36}{19^{2}}\approx 0.1.$
2.2 Sierpinski-like Triangles
While binary constructions at iterations $8k$, $k=1,2,\ldots$ reduce the figure to a spread of seeds, a distinct behavior emerges in a specific subsequence of iterations:
$i=2^{k+1}-1,\quad k=0,1,2,\ldots.$
At these steps, the figures reach their maximum local density (see Fig. 8) and, remarkably, form Sierpinski-like triangles—a behavior observed independently of the initial seed configuration.
An example of Sierpinski-like triangle growth is illustrated in Fig. 5(a). The estimated fractal dimension of the figure at the sixth iteration, obtained via the box-counting method, is approximately:
$D_{f}\approx 1.51.$
A broader classification of five types of Sierpinski-like figures, each arising from different seeds but following the same iterative process, is shown in Fig. 4.
Figure 4: Sierpinski-like triangle constructions obtained as $2$-style figures at iterations $i=2^{k+1}-1$, where $k=0,1,2,\ldots$. The same neighborhood applied to different seeds produces structurally similar results, though the fractal dimension (via box-counting) varies depending on the seed.
Additionally, we compare the $2$-style and 2322-style figures generated at the same subsequence of iterations in Fig. 5. Notably, while $2$-style figures form clear, structured Sierpinski-like patterns, the 2322-style construction introduces greater structural complexity, leading to an apparently more chaotic formation.
(a) Refer to caption
(b) Refer to caption
Figure 5: Comparison of $2$-style and 2322-style constructions from the same seeds and neighborhoods at the iteration sequence $i=2^{k+1}-1$, where $k=0,1,\ldots,5$. The 2322-style construction exhibits more irregularity and disorder compared to the structured Sierpinski-like growth of $2$-style figures.
3 Unique Properties of $2322$-Style Figures
The 2322-style construction, unlike purely binary ($2$) figures, introduces ternary interactions, resulting in distinctive structural and dynamical properties. These include:
  • 1.
    Lower density variance across iterations.
  • 2.
    Minimal connectivity loss (at most two steps away from full connectivity).
  • 3.
    Non-repetitive structural variation, distinct from the periodicity seen in $2$ figures.
At higher iterations, 2322-style figures retain complexity, while higher-order $222n$ 22-figures (for $n=5,7,9$) tend to exhibit more uniform growth patterns (Fig. 6).
(a)Refer to caption
(b)Refer to caption
[…]
3.1 Connectivity and Structural Complexity
The distinct connectivity properties of 2322-style figures can be observed in their early evolution. Fig. 7 illustrates how ternary interactions prevent the dissipation of seeds, which is characteristic of purely binary figures at iteration 8.
Figure 7: Comparison of $2$ and 2322 evolutions from the same seed. Unlike $2$ figures, 2322-style figures maintain structural connectivity, preventing dissociation into dispersed seeds at iteration 8.
[…]
  • 1.
    It is the most connected, requiring at most two steps to fully connect all components.
  • 2.
    Figures for $n=5,7,9$ tend to lose structural diversity, converging into similar shapes differing mainly in color distribution.
  • 3.
    2322 figures remain structurally distinct, retaining dynamic, non-repetitive features.
3.2 Density Variability and Aperiodic Behavior
The density evolution of $2$ and 2322 figures over time is compared in Fig. 8. Notably, $2$ figures undergo periodic density reductions at iterations $8k$, $k=1,2,\dots$, corresponding to their spread-of-seeds phenomenon. This is absent in 2322-style figures, which exhibit a more consistent density profile.
(a) Refer to caption
(b) Refer to caption
Figure 8: Comparison of densities of figures over iterations, Seed: One Point, Neighborhood: Diag Neumann.
(a) $2$-style and 2322-style figures over iterations. Local minima at iterations $8k$ correspond to seed dissociation in $2$ figures, a behavior not observed in 2322 figures.
[…]
Furthermore, we analyze density trends for odd-$n$ figures in Fig. 8(b). While figures for $n=5,7,9$ overlap, suggesting uniform periodicity, the 2322 figure remains structurally different.
3.3 Connection to Aperiodic Sequences
The non-periodic structural variation of 2322-style figures suggests connections to aperiodic mathematical structures, including:
  • 1.
    Dekking’s Construction: A non-repetitive sequence model in combinatorial mathematics.
  • 2.
    Quasi-periodic tilings: Found in quasicrystals and aperiodic lattices.
  • 3.
    Fractal substitution sequences, seen in hierarchical dynamical systems.
These findings reinforce the idea that modular arithmetic sequences within discrete Laplacians can yield quasi-periodic, self-organizing patterns, with potential applications in mathematical modeling, materials science, and biological growth simulations.
4 Statistical Comparison of $2$ and 2322 Sequences
This section investigates the numerical properties of discrete Laplacian sequences generated by $2$-style and 2322-style constructions at a fixed lattice cell. The goal is to determine whether these structures exhibit long-range correlations, spectral properties, and fractal characteristics, providing potential links to Dekking’s Construction and other non-periodic sequences.
4.1 Statistical Characteristics
Table 1 summarizes key statistical properties over 500 iterations. The entropy measures the degree of disorder, while variance and fractal dimension indicate complexity.
Sequence Type Entropy Mean Variance Fractal Dimension
$2$-style 0.1721 0.0275 0.0253 1.432
2322-style 1.0517 0.3318 0.3425 1.867
Table 1: Comparison of entropy, statistical measures, and fractal dimensions for $2$ and 2322 sequences over 500 iterations.
The higher entropy of the 2322 sequence indicates greater unpredictability compared to the periodic $2$ sequence. Similarly, the increased variance and higher fractal dimension suggest that 2322 structures exhibit greater structural complexity.
4.2 Fourier Spectral Analysis
Figures 9 and 10 show the Fourier transform of the sequences. The $2$ sequence exhibits sharp peaks, indicative of strong periodicity, whereas the 2322 sequence has a broad, distributed spectrum, suggesting quasi-periodicity or hierarchical self-organization.
Figure 9: Fourier spectrum of the $2$ sequence. Sharp peaks indicate strong periodicity.
Figure 10: Fourier spectrum of the 2322 sequence. The broader spectrum suggests quasi-periodicity or hierarchical structure.
The absence of clear periodic peaks in the 2322 case supports the hypothesis that it resembles non-periodic substitution sequences, such as those found in Dekking’s Construction. This aligns with known fractal systems and aperiodic tilings, where self-similar structures emerge without strict repetition.
4.3 Autocorrelation Analysis
The autocorrelation function, shown in Figures 11 and 12, provides further insight into sequence regularity.
  • 1.
    The $2$ sequence exhibits clear periodic correlations, confirming its structured, repeating nature—akin to the Sierpiński triangle.
  • 2.
    The 2322 sequence, in contrast, displays weaker long-range correlations and an irregular decay, suggesting a lack of strict periodicity.
Figure 11: Autocorrelation function of the $2$ sequence. Periodic correlations indicate a structured, repeating pattern.
Figure 12: Autocorrelation function of the 2322 sequence. The irregular decay suggests non-trivial long-range correlations.
This aperiodic structure suggests that 2322 figures belong to a distinct mathematical class, exhibiting properties found in quasicrystals, aperiodic tilings, and hierarchical dynamical systems.
4.4 Connection to Dekking’s Construction
Dekking’s Construction Dekking (1979) generates non-periodic sequences that avoid Abelian repetitions. The 2322 sequence, with its high entropy, irregular autocorrelation, and absence of Fourier periodicity, aligns closely with this framework.
  • 1.
    $2$ figures behave as classical fractals, with periodic behavior and predictable density fluctuations.
  • 2.
    2322 figures exhibit aperiodic, self-organizing growth, leading to a higher fractal dimension and more complex scaling properties.
These findings suggest that discrete Laplacians could serve as a graphical realization of Dekking-like sequences, providing new insights into hierarchical self-similarity, quasi-periodicity, and aperiodic mathematical structures.
4.5 Implications for Fractal and Aperiodic Systems
The observed aperiodicity and fractal-like behavior in 2322 sequences could have applications in:
  • 1.
    Quasicrystals and Aperiodic Order: The lack of strict periodicity is reminiscent of quasicrystals and other self-organizing materials.
  • 2.
    Computational Mathematics: Potential use in graph algorithms, hierarchical clustering, and AI-driven optimizations.
  • 3.
    Physical and Biological Systems: Non-repetitive structures appear in self-assembled molecular networks, biological growth patterns, and stochastic neural models.
The higher fractal dimension suggests that 2322-style figures do not conform to classical self-similar fractal models, reinforcing the idea that they belong to a distinct mathematical category, governed by non-trivial modular arithmetic and aperiodic order.
4.6 Future Research Directions
Future studies could explore:
  • 1.
    Higher-dimensional extensions: Extending discrete Laplacian iterations to 3D lattices to study their higher-rank algebraic properties.
  • 2.
    Stochastic variations: Investigating randomized Laplacian growth to model natural fractal phenomena.
  • 3.
    Mathematical characterization: Developing a rigorous classification of ternary interference in modular growth patterns.
These results suggest that the 2322 sequence represents a novel class of aperiodic, fractal-like structures, meriting further investigation in both pure mathematics and applied physics.
5 Conclusions and Applications
This study has explored the iterative construction of discrete Laplacians on 2D square lattices, focusing on the emergent properties of 2322-style patterns. Unlike $2$-style figures, which tend to display periodic behavior and seed dissociation at specific iterations, 2322-style figures exhibit quasi-aperiodicity, minimal connectivity loss, and low density variance. The self-similar yet non-repetitive nature of these figures suggests deep connections to aperiodic sequences, fractal geometry, and material organization.
Our findings have implications across mathematics, physics, biology, and computational science, reinforcing the role of discrete Laplacians as a framework for understanding complex hierarchical structures.
5.1 Mathematical Implications
  • 1.
    Aperiodic Sequences and Non-Repetitive Growth: The structural complexity of 2322-style figures aligns with Dekking’s Construction, which generates strongly non-repetitive sequences in combinatorial mathematics Dekking (1979). The alternating binary-ternary structure prevents periodic breakdown, making these figures distinct from traditional cellular automata.
  • 2.
    Fractal Properties: Box-counting methods estimate a fractal dimension of figures of approximately between 0.9 to 1.87, reinforcing the self-similar yet non-strictly repeating nature of these figures.
  • 3.
    Higher-Dimensional Extensions: Extending the iterative Laplacian process to 3D lattices could uncover higher-rank algebraic structures and new combinatorial sequences relevant to discrete geometry.
5.2 Physical and Material Science Applications
  • 1.
    Quasicrystals and Aperiodic Order: The quasi-periodicity of 2322-style figures is reminiscent of quasicrystalline structures, where long-range order exists without strict translational symmetry Gell-Mann (1994).
  • 2.
    Energy Dissipation in Materials: Valianti (2015) noted that modular arithmetic-based structures govern energy minimization and stochastic relaxation in complex alloys Valianti et al. (2015). The structured, yet non-repetitive, growth of 2322 figures suggests a possible role in fracture mechanics and self-healing materials.
  • 3.
    Fractal-Based Antennas, Sensors and Packages: Fractal-like structures have been successfully applied in sensor technology and electromagnetic wave manipulation Punjala and Makki (2009). The controlled irregularity of 2322 figures could inspire new designs for frequency-tuned fractal antennas and of properties of packages Zhang et al. (2021).
5.3 Biological and Computational Relevance
  • 1.
    Self-Organizing Fractals in Biology: Certain bacterial colonies and molecular self-assembly processes exhibit fractal-like growth patterns, with protein complexes in cyanobacteria forming Sierpiński-like triangles Sendker et al. (2024); Watson et al. (2023); Singh et al. (2024). This suggests that iterative Laplacians may serve as a mathematical model for biological self-organization.
  • 2.
    Neural Network Architectures and Structured Randomness: The non-uniform yet structured density of 2322 figures could inspire novel computational architectures, particularly in graph-based AI, sparse neural networks, and adaptive learning algorithms.
  • 3.
    Graph Algorithms and Pattern Recognition: The scaling properties of these figures could be leveraged in hierarchical clustering, AI-driven optimization, and modular computational frameworks.
5.4 Future Work
This study opens several research directions, including:
  • 1.
    Higher-dimensional discrete Laplacians: Investigating the properties of quaternary and quinary iterations in 3D and hypercubic lattices.
  • 2.
    Stochastic Variations and Randomized Growth Models: Introducing randomized Laplacian evolution to simulate biological fractals and chaotic systems.
  • 3.
    Experimental Applications in Material Science and Sensor Design: Exploring the real-world implementation of these patterns in metamaterials, photonic crystals, and electromagnetic wave applications.
By uncovering deep structural connections between discrete Laplacians, fractal dynamics, and aperiodic sequences, this work provides a unifying framework for mathematical exploration and applied sciences.
Appendix: Proof Sketch
We consider an initial seed contained in a $3\times 3$ square and apply the Diag-Neumann neighborhood, defined by:
$Nei=\begin{bmatrix}1&0&1\\ 0&0&0\\ 1&0&1\end{bmatrix}$
Let the initial seed $F_{0}$ be:
$F_{0}=\begin{bmatrix}a&b&c\\ d&\mathbf{e}&f\\ g&h&i\end{bmatrix}$
where $a,b,c,d,e,f,g,h,i\in\{0,1\}$ represent binary values.
We show that for $i=8k,k=1,2,3,\ldots$, the figure consists of a spread of its seeds and is at least 13-steps away from connectedness.
Step 1: First Iteration $F_{1}$
Applying the Laplacian update:
$a^{1}_{x,y}=\sum_{(u,v)\in Ne(x,y)}\left(a^{0}_{x+u,y+v}-a^{0}_{x,y}\right)\mod 2$
yields:
$F_{1}=\begin{bmatrix}a&b&ac&b&c\\ d&e&df&e&f\\ ag&bh&\mathbf{acgi}&bh&ci\\ d&e&df&e&f\\ g&h&gi&h&i\end{bmatrix}$
At this step, the pattern has expanded outward but remains connected.
Step 2: Second Iteration $F_{2}$
Continuing with the same update rule, we obtain:
$F_{2}=\begin{bmatrix}F_{0}&0_{3\times 1}&F_{0}\\ 0_{1\times 3}&0&0_{1\times 3}\\ F_{0}&0_{3\times 1}&F_{0}\end{bmatrix}$
At this stage, the pattern doubles in size but already shows a structure of dispersed seeds.
Step 3: Third Iteration $F_{3}$
Now we obtain:
$F_{3}=\begin{bmatrix}a&b&ac&b&ac&b&ac&b&c\\ d&e&df&e&df&e&df&e&f\\ ag&bh&acgi&bh&cdgi&bh&acgi&bh&ci\\ d&e&df&e&df&e&df&e&f\\ ag&bh&acgi&bh&cdgi&bh&acgi&bh&ci\\ d&e&df&e&df&e&df&e&f\\ g&h&gi&h&gi&h&gi&h&i\end{bmatrix}$
This is the last iteration before modular 2 cancellation starts to take effect.
Step 4: Fourth Iteration $F_{4}$
Applying the Laplacian update again, we obtain:
$F_{4}=\begin{bmatrix}F_{0}&0_{5\times 3}&F_{0}\\ 0_{5\times 3}&0_{5\times 5}&0_{5\times 3}\\ F_{0}&0_{5\times 3}&F_{0}\end{bmatrix}$
Steps 5-7: Expansion of Each $F_{0}$ Within $F_{4}$
Each of the four $F_{0}$ components in $F_{4}$ will now expand similarly to $F_{3}$ at the seventh step. That means we are one step away from a fully expanded spread.
Step 8: Modular 2 Cancellation
At iteration $i=8$, the only additional effect is that overlapping rows and columns will contribute extra values. However, because of mod 2 arithmetic, any double additions cancel out to 0.
Thus, the structure at step $F_{8}$ consists of four separate instances of $F_{0}$ spread out in a larger matrix, with gaps remaining between them. This completes the sketch of the proof. $\square$
Declarations
  • 1.
    Funding: Not applicable.
  • 2.
    Conflicts of Interest: The author declares no conflicts of interest.
  • 3.
    Author Contributions: Małgorzata Nowak-Kȩpczyk conducted all research and analysis.

### v2

@@
 We investigate the iterative construction of discrete Laplacians on 2D square lattices, revealing emergent fractal-like patterns shaped by modular arithmetic.
-While classical $2$-style iterations reproduce known structures such as the Sierpiński triangle, our alternating binary–ternary (2322-style) process produces a novel class of aperiodic figures.
+While classical 2222-style iterations reproduce known structures such as the Sierpiński triangle, our alternating binary–ternary (2322-style) process produces a novel class of aperiodic figures.
 These display low density variance, minimal connectivity loss, and non-repetitive organization reminiscent of Dekking’s sequences.
-Fourier and autocorrelation analyses confirm their quasi-periodic nature, suggesting applications in self-assembly, sensor networks, and biological modeling.
-The findings open new paths toward structured randomness and fractal dynamics in discrete systems.
-These findings also open avenues for exploring higher-dimensional Laplacian constructions and their implications in quasicrystals, aperiodic tilings, and stochastic processes.
+Fourier and autocorrelation analyses confirm their quasi-periodic nature, suggesting structural analogies with self-assembly processes and fractal-based design.
+The results highlight the structural richness of modular Laplacian dynamics and motivate further study of non-periodic discrete systems.
 keywords:
@@
 Iterating the discrete Laplacian on 2D square lattices gives rise to complex emergent structures, often displaying fractal-like organization.
-Previous studies of Discrete Laplacians on 2D lattices have documented visually striking fractal patterns, such as snowflake-like and carpet-like motifs, arising from binary iterations alone Aiba et al.
-(2006).
+Previous studies of Discrete Laplacians on 2D lattices have documented visually striking fractal patterns, such as snowflake-like and carpet-like motifs, arising from binary iterations alone [1].
 These patterns emerge due to modular arithmetic constraints and exhibit diverse forms, depending on the chosen seed and neighborhood structure.
@@
 This motivates the exploration of higher-order modular arithmetic, such as ternary (modulo 3) and quaternary (modulo 4) iterations, which significantly expand the space of possible patterns and often result in aperiodic, structurally richer formations.
+Earlier investigations of modular Laplacian dynamics in the purely binary setting are presented in [5].
 In this paper, we investigate a special class of iterative Laplacian constructions, focusing on the interplay between binary and ternary arithmetic in pattern formation.
 We introduce the 2322-style pattern, an alternating sequence of binary and ternary iterations, which gives rise to structurally distinct, non-repetitive figures.
-Unlike purely binary $2$-style figures, which exhibit dissociative periodicity, 2322-style figures display low density variance, minimal connectivity loss, and quasi-aperiodicity.
+Unlike purely binary 2222-style figures, which exhibit dissociative periodicity, 2322-style figures display low density variance, minimal connectivity loss, and quasi-aperiodicity.
 These properties are reminiscent of Dekking’s non-repetitive sequences, which play a significant role in combinatorial mathematics and aperiodic order.
-Beyond their mathematical significance, these structures may have applications in physics, materials science, and computational modeling.
+Beyond their mathematical significance, these structures may have broader structural implications.
 The concluding section of this paper will discuss their potential relevance, including connections to biological self-organization, sensor technology, material science, and fractal-based computational algorithms.
@@
 1.2 Automaton Evolution
-The iterative update rule follows the form:
-$\nu_{i}(p)=\sum_{g\in Ne(p)}\big{(}\nu(g)-\nu(p)\big{)}\mod n,$ (1)
-where:
+Let $G=(V,E)$ denote the square lattice graph, where $V$ is the set of lattice points and $E$ is determined by the chosen neighborhood structure. For $p\in V$, we denote by
+$Ne(p)=\{g\in V:(p,g)\in E\}$
+the set of neighboring vertices of $p$.
+Consider a sequence
+$\mathcal{S}=(k_{1},k_{2},k_{3},\ldots),\qquad k_{i}\in\{2,3,4,\ldots\}.$
+The discrete graph Laplacian applied to $\nu_{i}$ is defined by
+$(\Delta\nu_{i})(p)=\sum_{g\in Ne(p)}\big(\nu_{i}(g)-\nu_{i}(p)\big)$ (1)
+defined over $\mathbb{Z}$ prior to modular reduction. The evolution rule is given by
+$\nu_{i+1}=(\Delta\nu_{i})\mod k_{i},$
+where $k_{i}$ is the $i$-th element of $\mathcal{S}$.
+Here:
   • 1.
-    $n$ is the modulus (defining binary, ternary, or higher-order arithmetic),
+    $k_{i}$ is the modulus (binary for $k_{i}=2$, ternary for $k_{i}=3$, etc.),
   • 2.
-    $i$ represents the iteration step,
+    $i$ denotes the discrete time (iteration) step,
   • 3.
-    $Ne(p)$ defines the neighborhood structure, governing local interactions.
+    $Ne(p)$ defines the neighborhood structure (e.g., von Neumann or diagonal).
+Thus, the automaton evolves by repeated application of the graph Laplacian, followed by reduction modulo $k_{i}$, for $i=1,2,3,\ldots$.
 Figure 2: Examples of neighborhoods.
 1.3 Modular Arithmetic and Iterative Extensions
-To extend beyond binary figures, we introduce modular arithmetic with $n>2$. Each color represents a distinct modular class:
+The introduction of the sequence $\mathcal{S}$ allows us to extend the construction beyond binary (black-and-white) figures, which correspond to the case when $\mathcal{S}$ is constant, i.e.
+$\mathcal{S}=(2,2,2,\ldots)$.
+In particular, for constant sequences $\mathcal{S}=(n,n,n,\ldots)$ we obtain:
   • 1.
-    Binary iteration ($n=2$): classical black-and-white figures.
+    Ternary iteration ($n=3$): a three-state structure with increased combinatorial complexity.
   • 2.
-    Ternary iteration ($n=3$): introduces a new state, increasing complexity.
-  • 3.
-    Quaternary and higher ($n>3$): potential multi-state structures.
-We define a $k$-nary iteration:
-$u_{i}(p)=\sum_{g\in Ne(p)}\big{(}u(g)-u(p)\big{)}\mod k,\quad k=2,3,4,\dots$
-where the iteration is binary for $k=2$, ternary for $k=3$, and so forth.
-In this paper, we examine specific iteration sequences:
-$u_{i}(p)=\begin{cases}\sum\limits_{g\in Ne(p)}\big{(}u(g)-u(p)\big{)}\mod n,&\text{if }i\equiv 2\pmod{4},\\ \sum\limits_{g\in Ne(p)}\big{(}u(g)-u(p)\big{)}\mod 2,&\text{otherwise}.\end{cases}$ (2)
-This generates $222n$ 22-style figures, including:
+    Quaternary and higher ($n>3$): multi-state structures with $n$ distinct states.
+In this paper, we focus on sequences $\mathcal{S}$ of the form
+$k_{i}=\begin{cases}n,&\text{if }i\equiv 2\pmod{4},\\ 2,&\text{otherwise},\end{cases}$ (2)
+and the corresponding family of figures generated by the Laplacian dynamics will be referred to as $2n22$-style figures [^1] Other families of construction sequences are studied in [6].
+In particular:
   • 1.
-    $2$-style (purely binary evolution),
+    2222-style: purely binary evolution ($n=2$),
   • 2.
-    2322-style (introducing ternary interference in specific steps).
-Figure 3: Binary iterations from a Sierpiński seed and Neumann neighborhood.
+    2322-style: ternary interference at every fourth step ($n=3$).
+Figure 3: Binary ($2222$-style) evolution from a three-point seed under the von Neumann neighborhood.
+The upper left panel shows the initial seed and the assumed neighborhood structure.
+The subsequent panels display the 0-th (initial configuration), first, second, and third iterations generated by the modular Laplacian dynamics with $k_{i}=2$, illustrating the initial expansion induced by the discrete Laplacian.
 2 Binary style figures and their properties
-Prior work on binary style figures by Suzuki and Maegaito uncovered visually striking configurations, including snowflake-like, butterfly-like, and Persian carpet-like motifs Aiba et al.
-(2006).
-Theoretical developments, including fixed-point theorems, periodicity results, and connections to binomial and trinomial sequences have been found Hadlich et al.
-(2011).
+Prior work on binary style figures by Suzuki and Maegaito uncovered visually striking configurations, including snowflake-like, butterfly-like, and Persian carpet-like motifs [1].
+Theoretical developments, including fixed-point theorems, periodicity results, and connections to binomial and trinomial sequences have been found [4].
 We shall investigate binary style figures via many seeds and neighbourhoods. First we shall describe characteristic feature of binary figures and the suprising appearance of Sierpinski-like triangles.
@@
 A connected figure ($A$), figures: 1-step away (figure $B$), and 2-steps away from connectedness (figure $C$) are shown below.
-All binary-style figures posses a special feature which is described in:
-Proposition. Every binary-style figure of arbitrary seed and neighborhood at iterations $i=8k$, $k=1,2,3,\ldots$ consists of a spread of its seeds and is at least 13-steps away from connectedness.
+Proposition.
+Let the binary Laplacian dynamics be defined with the Diag-Neumann neighborhood and let the initial configuration be finitely supported.
+Then for every $k\geq 1$, the configuration at iteration $i=8k$ consists of disjoint translated copies of the initial seed.
+Moreover, these copies are separated by a positive gap (which equals $13$ in the $3\times 3$ case considered here).
 An illustration of this proposition is shown in Fig. 3. The sketch of the proof will be given in the Appendix.
-Generalization to Other Neighborhoods
-Although this proof is carried out for Diag-Neumann, all observed figures for various seeds and neighborhoods exhibited the same behavior.
-Therefore, it is highly likely that all $222n$ 22-style figures for even $n$ behave similarly, all the figures are reduced to a spread of seeds at iterations $8k$.
+Remark on other neighborhoods.
+The above argument is carried out for the Diag-Neumann neighborhood.
+Numerical experiments indicate that similar replication phenomena occur for other neighborhood structures as well.
+A general operator-based proof for arbitrary neighborhoods is beyond the scope of the present paper.
 2.1 Density Measure
@@
 Definition. The density $\rho(i)$ of a figure at iteration $i$ is defined as the ratio of occupied cells to the total available lattice area:
-$\rho(i)=\frac{\sum_{p\in\text{lattice}}\text{sgn}(f(p))}{(3+2i)^{2}}.$
+$\rho(i)=\frac{\sum_{p\in\text{lattice}}\text{sgn}(\nu(p))}{(3+2i)^{2}}.$
 Since binary figures are observed to become a spread of seeds at iterations $i=8k,\,k=1,2,\dots$, their density exhibits sharp reductions at these steps (see Fig.
@@
 A broader classification of five types of Sierpinski-like figures, each arising from different seeds but following the same iterative process, is shown in Fig. 4.
-Figure 4: Sierpinski-like triangle constructions obtained as $2$-style figures at iterations $i=2^{k+1}-1$, where $k=0,1,2,\ldots$.
+Figure 4: Sierpinski-like triangle constructions obtained as 2222-style figures at iterations $i=2^{k+1}-1$, where $k=0,1,2,\ldots$.
 The same neighborhood applied to different seeds produces structurally similar results, though the fractal dimension (via box-counting) varies depending on the seed.
-Additionally, we compare the $2$-style and 2322-style figures generated at the same subsequence of iterations in Fig.
+Additionally, we compare the 2222-style and 2322-style figures generated at the same subsequence of iterations in Fig.
 5.
-Notably, while $2$-style figures form clear, structured Sierpinski-like patterns, the 2322-style construction introduces greater structural complexity, leading to an apparently more chaotic formation.
+Notably, while 2222-style figures form clear, structured Sierpinski-like patterns, the 2322-style construction introduces greater structural complexity, leading to an apparently more chaotic formation.
 (a) Refer to caption
 (b) Refer to caption
-Figure 5: Comparison of $2$-style and 2322-style constructions from the same seeds and neighborhoods at the iteration sequence $i=2^{k+1}-1$, where $k=0,1,\ldots,5$.
-The 2322-style construction exhibits more irregularity and disorder compared to the structured Sierpinski-like growth of $2$-style figures.
+Figure 5: Comparison of 2222-style and 2322-style constructions from the same seeds and neighborhoods at the iteration sequence $i=2^{k+1}-1$, where $k=0,1,\ldots,5$.
+The 2322-style construction exhibits more irregularity and disorder compared to the structured Sierpinski-like growth of $22$-style figures.
 3 Unique Properties of $2322$-Style Figures
-The 2322-style construction, unlike purely binary ($2$) figures, introduces ternary interactions, resulting in distinctive structural and dynamical properties. These include:
+The 2322-style construction, unlike purely binary (2222) figures, introduces ternary interactions, resulting in distinctive structural and dynamical properties. These include:
   • 1.
@@
   • 3.
-    Non-repetitive structural variation, distinct from the periodicity seen in $2$ figures.
-At higher iterations, 2322-style figures retain complexity, while higher-order $222n$ 22-figures (for $n=5,7,9$) tend to exhibit more uniform growth patterns (Fig. 6).
+    Non-repetitive structural variation, distinct from the periodicity seen in 2222 figures.
+At higher iterations, 2322-style figures retain complexity, while higher-order $2n22$-figures (for $n=5,7,9$) tend to exhibit more uniform growth patterns (Fig. 6).
 (a)Refer to caption
 (b)Refer to caption
-[…]
-[…]
+Figure 6: Seed: single point; diagonal Neumann neighborhood.
+(a) Comparison of $2n22$ figures for $n=2,3,4,5$ (left to right; iterations 30–32,34).
+The 2322 case (second column) is structurally distinct.
+Fractal dimensions for 2222 (resp.
+2322) are 1.31, 1.66, 0.91, 1.14 (resp.
+1.78, 1.62, 1.43, 1.62).
+(b) Comparison of 2322 with $2n22$ for odd $n=3,5,7,9$ (iterations 23–25,27), showing persistent structural distinctness.
 3.1 Connectivity and Structural Complexity
@@
 7 illustrates how ternary interactions prevent the dissipation of seeds, which is characteristic of purely binary figures at iteration 8.
-Figure 7: Comparison of $2$ and 2322 evolutions from the same seed.
-Unlike $2$ figures, 2322-style figures maintain structural connectivity, preventing dissociation into dispersed seeds at iteration 8.
-[…]
+Figure 7: Comparison of 2222 and 2322 evolutions from the same seed.
+Unlike 2222 figures, 2322-style figures maintain structural connectivity, in contrast to the seed dissociation observed in 2222 at 8ii-ths iterations.
+Among 2nn22-style figures for $n$ odd, 2322 exhibits distinct structural behaviour:
   • 1.
@@
 3.2 Density Variability and Aperiodic Behavior
-The density evolution of $2$ and 2322 figures over time is compared in Fig.
+The density evolution of 2222 and 2322 figures over time is compared in Fig.
 8.
-Notably, $2$ figures undergo periodic density reductions at iterations $8k$, $k=1,2,\dots$, corresponding to their spread-of-seeds phenomenon.
+Notably, 2222 figures undergo periodic density reductions at iterations $8k$, $k=1,2,\dots$, corresponding to their spread-of-seeds phenomenon.
 This is absent in 2322-style figures, which exhibit a more consistent density profile.
@@
 Figure 8: Comparison of densities of figures over iterations, Seed: One Point, Neighborhood: Diag Neumann.
-(a) $2$-style and 2322-style figures over iterations. Local minima at iterations $8k$ correspond to seed dissociation in $2$ figures, a behavior not observed in 2322 figures.
-[…]
+(a) 2222-style and 2322-style figures over iterations. Local minima at iterations $8k$ correspond to seed dissociation in 2222 figures, a behavior not observed in 2322 figures.
+(b) 2nn22-style figures, $n=3,5,7,9$. Graphs for $n=5,7,9$ overlap, suggesting similar periodic behavior, whereas 2322 maintains a unique profile.
 Furthermore, we analyze density trends for odd-$n$ figures in Fig. 8(b). While figures for $n=5,7,9$ overlap, suggesting uniform periodicity, the 2322 figure remains structurally different.
@@
 These findings reinforce the idea that modular arithmetic sequences within discrete Laplacians can yield quasi-periodic, self-organizing patterns, with potential applications in mathematical modeling, materials science, and biological growth simulations.
-4 Statistical Comparison of $2$ and 2322 Sequences
-This section investigates the numerical properties of discrete Laplacian sequences generated by $2$-style and 2322-style constructions at a fixed lattice cell.
+4 Statistical Comparison of 2222 and 2322 Sequences
+This section investigates the numerical properties of discrete Laplacian sequences generated by 2222-style and 2322-style constructions at a fixed lattice cell.
 The goal is to determine whether these structures exhibit long-range correlations, spectral properties, and fractal characteristics, providing potential links to Dekking’s Construction and other non-periodic sequences.
 4.1 Statistical Characteristics
+The entropy is computed in the Shannon sense,
+$H=-\sum_{j}p_{j}\log p_{j},$
+where $p_{j}$ denotes the empirical frequency of the $j$-th state. The fractal dimension is estimated using a standard box-counting method.
 Table 1 summarizes key statistical properties over 500 iterations. The entropy measures the degree of disorder, while variance and fractal dimension indicate complexity.
 Sequence Type Entropy Mean Variance Fractal Dimension
-$2$-style 0.1721 0.0275 0.0253 1.432
+$22$-style 0.1721 0.0275 0.0253 1.432
 2322-style 1.0517 0.3318 0.3425 1.867
-Table 1: Comparison of entropy, statistical measures, and fractal dimensions for $2$ and 2322 sequences over 500 iterations.
-The higher entropy of the 2322 sequence indicates greater unpredictability compared to the periodic $2$ sequence.
+Table 1: Comparison of entropy, statistical measures, and fractal dimensions for 2222 and 2322 sequences over 500 iterations.
+The higher entropy of the 2322 sequence indicates greater unpredictability compared to the periodic 2222 sequence.
 Similarly, the increased variance and higher fractal dimension suggest that 2322 structures exhibit greater structural complexity.
 4.2 Fourier Spectral Analysis
-Figures 9 and 10 show the Fourier transform of the sequences.
-The $2$ sequence exhibits sharp peaks, indicative of strong periodicity, whereas the 2322 sequence has a broad, distributed spectrum, suggesting quasi-periodicity or hierarchical self-organization.
-Figure 9: Fourier spectrum of the $2$ sequence. Sharp peaks indicate strong periodicity.
+Figures 9 and 10 present the amplitude spectra of the corresponding sequences.
+Figure 9: Fourier spectrum of the 2222 sequence. Sharp peaks indicate strong periodicity.
 Figure 10: Fourier spectrum of the 2322 sequence. The broader spectrum suggests quasi-periodicity or hierarchical structure.
-The absence of clear periodic peaks in the 2322 case supports the hypothesis that it resembles non-periodic substitution sequences, such as those found in Dekking’s Construction.
-This aligns with known fractal systems and aperiodic tilings, where self-similar structures emerge without strict repetition.
+The 2222 sequence exhibits pronounced discrete peaks located at regularly spaced frequencies, indicating strong periodic components and a well-defined fundamental frequency.
+In contrast, the 2322 sequence does not display a single dominant periodic frequency.
+Instead, its spectrum consists of multiple distributed peaks with irregular spacing and varying amplitude, reflecting hierarchical spatial organization rather than strict periodic repetition.
+Thus, while the 2222 case corresponds to a predominantly discrete spectrum, the 2322 sequence exhibits a more distributed spectral structure, consistent with non-periodic yet structured growth.
 4.3 Autocorrelation Analysis
@@
   • 1.
-    The $2$ sequence exhibits clear periodic correlations, confirming its structured, repeating nature—akin to the Sierpiński triangle.
+    The 2222 sequence exhibits clear periodic correlations, confirming its structured, repeating nature—akin to the Sierpiński triangle.
   • 2.
     The 2322 sequence, in contrast, displays weaker long-range correlations and an irregular decay, suggesting a lack of strict periodicity.
-Figure 11: Autocorrelation function of the $2$ sequence. Periodic correlations indicate a structured, repeating pattern.
+This aperiodic structure suggests that 2322 figures belong to a distinct mathematical class, exhibiting properties found in quasicrystals, aperiodic tilings, and hierarchical dynamical systems.
+Figure 11: Autocorrelation function of the 2222 sequence. Periodic correlations indicate a structured, repeating pattern.
 Figure 12: Autocorrelation function of the 2322 sequence. The irregular decay suggests non-trivial long-range correlations.
-This aperiodic structure suggests that 2322 figures belong to a distinct mathematical class, exhibiting properties found in quasicrystals, aperiodic tilings, and hierarchical dynamical systems.
 4.4 Connection to Dekking’s Construction
-Dekking’s Construction Dekking (1979) generates non-periodic sequences that avoid Abelian repetitions.
-The 2322 sequence, with its high entropy, irregular autocorrelation, and absence of Fourier periodicity, aligns closely with this framework.
+Dekking’s Construction [2] generates non-periodic sequences that avoid Abelian repetitions.
+The 2322 sequence, with its high entropy, irregular autocorrelation, and absence of Fourier periodicity, shows structural similarities to this framework.
   • 1.
-    $2$ figures behave as classical fractals, with periodic behavior and predictable density fluctuations.
+    2222 figures behave as classical fractals, with periodic behavior and predictable density fluctuations.
   • 2.
@@
 These findings suggest that discrete Laplacians could serve as a graphical realization of Dekking-like sequences, providing new insights into hierarchical self-similarity, quasi-periodicity, and aperiodic mathematical structures.
-4.5 Implications for Fractal and Aperiodic Systems
-The observed aperiodicity and fractal-like behavior in 2322 sequences could have applications in:
-  • 1.
-    Quasicrystals and Aperiodic Order: The lack of strict periodicity is reminiscent of quasicrystals and other self-organizing materials.
-  • 2.
-    Computational Mathematics: Potential use in graph algorithms, hierarchical clustering, and AI-driven optimizations.
-  • 3.
-    Physical and Biological Systems: Non-repetitive structures appear in self-assembled molecular networks, biological growth patterns, and stochastic neural models.
-The higher fractal dimension suggests that 2322-style figures do not conform to classical self-similar fractal models, reinforcing the idea that they belong to a distinct mathematical category, governed by non-trivial modular arithmetic and aperiodic order.
 4.5 Future Research Directions
@@
 These results suggest that the 2322 sequence represents a novel class of aperiodic, fractal-like structures, meriting further investigation in both pure mathematics and applied physics.
-5 Conclusions and Applications
-This study has explored the iterative construction of discrete Laplacians on 2D square lattices, focusing on the emergent properties of 2322-style patterns.
-Unlike $2$-style figures, which tend to display periodic behavior and seed dissociation at specific iterations, 2322-style figures exhibit quasi-aperiodicity, minimal connectivity loss, and low density variance.
-The self-similar yet non-repetitive nature of these figures suggests deep connections to aperiodic sequences, fractal geometry, and material organization.
-Our findings have implications across mathematics, physics, biology, and computational science, reinforcing the role of discrete Laplacians as a framework for understanding complex hierarchical structures.
+5 Conclusions and Perspectives
+We have studied modular Laplacian dynamics on two-dimensional square lattices, with particular emphasis on non-constant sequences such as the $2322$ pattern.
+In contrast to purely binary ($2222$) constructions, which exhibit periodic behavior and seed dissociation at specific iterations, the $2322$-style figures display reduced periodicity, higher entropy, and increased fractal dimension.
+These results indicate that introducing variable modular arithmetic significantly modifies the long-term structural behavior of discrete Laplacian iterations.
 5.1 Mathematical Implications
   • 1.
-    Aperiodic Sequences and Non-Repetitive Growth: The structural complexity of 2322-style figures aligns with Dekking’s Construction, which generates strongly non-repetitive sequences in combinatorial mathematics Dekking (1979).
-The alternating binary-ternary structure prevents periodic breakdown, making these figures distinct from traditional cellular automata.
+    Non-periodic growth: The structural complexity of $2322$-style figures suggests analogies with non-periodic substitution systems, including constructions related to Dekking-type sequences [2].
+This correspondence is structural rather than formal.
   • 2.
-    Fractal Properties: Box-counting methods estimate a fractal dimension of figures of approximately between 0.9 to 1.87, reinforcing the self-similar yet non-strictly repeating nature of these figures.
+    Fractal characteristics: Box-counting estimates indicate fractal dimensions in the approximate range 0.90.91.871.87, consistent with the observed multi-scale self-similarity.
   • 3.
-    Higher-Dimensional Extensions: Extending the iterative Laplacian process to 3D lattices could uncover higher-rank algebraic structures and new combinatorial sequences relevant to discrete geometry.
-5.2 Physical and Material Science Applications
+    Higher-dimensional extensions: Extending modular Laplacian dynamics to three-dimensional lattices and other graph structures remains a natural direction for further study.
+5.2 Physical and Computational Perspectives
   • 1.
-    Quasicrystals and Aperiodic Order: The quasi-periodicity of 2322-style figures is reminiscent of quasicrystalline structures, where long-range order exists without strict translational symmetry Gell-Mann (1994).
+    Aperiodic order: The reduced periodicity of $2322$-style figures shows structural similarities to systems exhibiting aperiodic order, such as quasicrystalline patterns [3].
+This analogy concerns geometric organization rather than a direct physical model.
   • 2.
-    Energy Dissipation in Materials: Valianti (2015) noted that modular arithmetic-based structures govern energy minimization and stochastic relaxation in complex alloys Valianti et al.
-(2015).
-The structured, yet non-repetitive, growth of 2322 figures suggests a possible role in fracture mechanics and self-healing materials.
+    Fractal-inspired design: Self-similar geometries are widely used in electromagnetic and sensor applications [7, 11].
+The controlled irregularity of the $2322$-style figures suggests potential relevance for scalable pattern generation.
   • 3.
-    Fractal-Based Antennas, Sensors and Packages: Fractal-like structures have been successfully applied in sensor technology and electromagnetic wave manipulation Punjala and Makki (2009).
-The controlled irregularity of 2322 figures could inspire new designs for frequency-tuned fractal antennas and of properties of packages Zhang et al.
-(2021).
-5.3 Biological and Computational Relevance
+    Structured discrete models: The non-uniform yet organized density patterns provide an abstract framework for studying hierarchical structures in graph-based systems.
+  • 4.
+    Motivation from natural self-assembly: Recent observations of Sierpiński-like protein structures in biological systems [8, 10, 9] demonstrate that fractal growth patterns may arise in natural self-assembly processes.
+While the present work is purely mathematical, such findings provide motivation for studying discrete self-similar dynamics.
+5.3 Future Directions
+Possible directions for further investigation include:
   • 1.
-    Self-Organizing Fractals in Biology: Certain bacterial colonies and molecular self-assembly processes exhibit fractal-like growth patterns, with protein complexes in cyanobacteria forming Sierpiński-like triangles Sendker et al.
-(2024); Watson et al.
-(2023); Singh et al.
-(2024).
-This suggests that iterative Laplacians may serve as a mathematical model for biological self-organization.
+    Analytical characterization of non-binary modular sequences;
   • 2.
-    Neural Network Architectures and Structured Randomness: The non-uniform yet structured density of 2322 figures could inspire novel computational architectures, particularly in graph-based AI, sparse neural networks, and adaptive learning algorithms.
+    Stochastic variants and perturbed modular iterations;
   • 3.
-    Graph Algorithms and Pattern Recognition: The scaling properties of these figures could be leveraged in hierarchical clustering, AI-driven optimization, and modular computational frameworks.
-5.4 Future Work
-This study opens several research directions, including:
-  • 1.
-    Higher-dimensional discrete Laplacians: Investigating the properties of quaternary and quinary iterations in 3D and hypercubic lattices.
-  • 2.
-    Stochastic Variations and Randomized Growth Models: Introducing randomized Laplacian evolution to simulate biological fractals and chaotic systems.
-  • 3.
-    Experimental Applications in Material Science and Sensor Design: Exploring the real-world implementation of these patterns in metamaterials, photonic crystals, and electromagnetic wave applications.
-By uncovering deep structural connections between discrete Laplacians, fractal dynamics, and aperiodic sequences, this work provides a unifying framework for mathematical exploration and applied sciences.
+    Extensions to higher-dimensional lattices.
+These directions aim to deepen the mathematical understanding of modular Laplacian dynamics and their structural properties.
 Appendix: Proof Sketch
@@
 Step 8: Modular 2 Cancellation
-At iteration $i=8$, the only additional effect is that overlapping rows and columns will contribute extra values. However, because of mod 2 arithmetic, any double additions cancel out to 0.
-Thus, the structure at step $F_{8}$ consists of four separate instances of $F_{0}$ spread out in a larger matrix, with gaps remaining between them. This completes the sketch of the proof. $\square$
+At iteration $i=8$, the only additional effect comes from overlaps between the expanding translated blocks.
+However, since the evolution is performed modulo $2$, any contributions appearing twice cancel out.
+The translation offset at this stage equals $16$ in each coordinate direction.
+Since the initial seed is contained in a $3\times 3$ square, its maximal diameter does not exceed $3$.
+Therefore, the minimal separation between the translated copies at iteration $8$ is given by
+163=13.16-3=13.
+Thus, the configuration $F_{8}$ consists of four disjoint translated copies of the initial seed, separated by a positive gap. This completes the sketch of the proof. $\qed$
 Declarations